OpenStudy (anonymous):

Let f(x)=-2x-9 with domain I. Find the following. a.f(-7) b.f(100) c.f(-a)in terms of a

OpenStudy (campbell_st):

substitute each value seperately... x = -7, x = 100 then evaluate. for x = -a just substitute and evaluate.... but you will still end up with an equation

OpenStudy (anonymous):

not understanding, a? b? c?

OpenStudy (campbell_st):

ok...for the 1st part find \[ f(-7) = -2 \times -7 - 9\] you need to evaluate... get a number answer for this

OpenStudy (anonymous):

Please help me, I am not understanding any of this

OpenStudy (anonymous):

Ok, so I got 5 for the first, now the second?

OpenStudy (campbell_st):

ok... here it is again \[f(-7) = ( - 2 \times -7) - 9\] can you get a number answer for the right hand side of the equation...?

OpenStudy (anonymous):

I got 5 for a, what about b and c??

OpenStudy (campbell_st):

ok... thats the 1st part... repeat the above for x = 100 so in the above equation replace - 7 with 100 then for part c replace - 7 with - a

OpenStudy (anonymous):

-209 for b

OpenStudy (campbell_st):

thats correct...

OpenStudy (anonymous):

Ok so what is the anwer for c???

OpenStudy (campbell_st):

\[f(-a) = (-2 \times -a) - 9\] simplify it... and it will be an equation

OpenStudy (anonymous):

Is it -2x-a-9?

OpenStudy (campbell_st):

no remember yo are replacing x with -a

OpenStudy (campbell_st):

and in the equation I've written its a multiplication sign between -2 and -a

OpenStudy (anonymous):

I don't know, I keep getting the anwswer wrong, I only have one more try to get the answer correct, please help me

OpenStudy (campbell_st):

what is -2 times -a ...?

OpenStudy (anonymous):

2a?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok, so what is the answer for c?

OpenStudy (campbell_st):

correct... so f(-a) = 2a - 9

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Ok, I got it!

OpenStudy (campbell_st):

hope it helps

OpenStudy (anonymous):

Yes, thank you so much

OpenStudy (anonymous):

I have one more problem

OpenStudy (campbell_st):

sorry... I have to go... but there are lots of people here who can help... good luck